Joe's Dairy Farm has 100 gallons of milk that is 4.6% butterfat. How much skimmilk (no butterfat) should be mixed with it to make milk that is 3.2% fat?
________gallons of milk should be added.
--------------------------------------------------------------------------------
A chemist has one solution that is 80% acid and another soulution that is 30% acid. How much of the first (80%) solution is needed to make a 400L solution that is 62% acid?
__________liters of the 80% solution is required.
Thanx

Answer by longjonsilver(2297)

read my Lesson on Mixtures. That should help you understand this type of question.

Anyway:
1. We cannot think about "amounts of butterfat, since the skimmed has none: its percentage is zero. So i have turned it around and considered the percentage of non-fat milk in the milk: the skimmed will be 100%.

100*0.954 + x*1.00 = (100+x)*0.968
95.4 + x = 96.8 + 0.968x
x - 0.968x = 96.8 - 95.4
0.032x = 1.4
x = (1.4)/(0.032)
x = 43.75 gallons

2. Start with 80% acid and a 30% acid.
--> Let there be x litres of the 80% solution acid. So, how much of the 30%? Answer is "the rest of the 400 litres ie (400-x). This is the tricky bit you need to "see".

So, 0.80x + 0.30(400-x) = 0.62*400
0.80x + 120 - 0.30x = 256
0.80x - 0.30x = 256 - 120
0.50x = 136
x = 272

So we need 272 litres of the 80% acid
and we need 128 litres of the 30% acid

jon.